OpenStudy (anonymous):

Use the value of the trigonometric function to evaluate the indicated function. sin t=1/2 csc (-t)=?

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ csc(\theta)=\cfrac{1}{sin(\theta)}\qquad sin(-\theta)=-sin(\theta)}} \\ \quad \\ sin(t)=\cfrac{1}{2}\qquad csc(-t) = \cfrac{1}{sin(-t)}\implies ?\)